OpenStudy (anonymous):

x^3+y^3=6xy (4/3, 8/3) find slope

OpenStudy (anonymous):

Take the derivative, That makes find the slope very simple

OpenStudy (anonymous):

Are you asked to find the slope at that point? BTW, do you know how to take a derivative or a polynomial?

OpenStudy (anonymous):

I don't really understand derivatives. That's where the problem starts

OpenStudy (anonymous):

The slope at that point

OpenStudy (anonymous):

I'll take the derivative for you, could you double check your equation?

OpenStudy (anonymous):

x^3+y^3=6xy

OpenStudy (anonymous):

alright, the derivative is (x^2-2y)/(2x-y^2), now just plug in the point you have and that will give you the slope of the lline at that point.

OpenStudy (anonymous):

I'm getting the slope to be 4/5

OpenStudy (anonymous):

How do you get to the derivative though?

OpenStudy (anonymous):

What math class/ level are you in?

OpenStudy (anonymous):

You just do implicit differentiation. This is because y is part of the equation and it can't be isolated.

OpenStudy (anonymous):

high school calculus

OpenStudy (anonymous):

Have you done derivatives, or implicit differentiation?

OpenStudy (anonymous):

working n implicit

OpenStudy (anonymous):

Do you know how to start? Take the derivative dy/dx(x^3+y^3)=d/dx(6xy)